# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v7

@@ Abstract
Objective: The proposed method deciphers electrical brain states which as multi-dimensional space-neutral cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
@@ 1 Introduction
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel space-neutral method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
@@ 2.3 Brain states detection (BSD)
In a first move a limited set of eight brain states are computed. Raw EEG signals from each channel are transformed into sequences of overlapping spectrograms using the Fast Fourier Transform (FFT) implementation from the cusignal.spectrogram library (Thompson and Nicely, 2021). The recordings are segmented into data epochs where each individual window, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz). While window size and number of states are varied as technical optimization against detection performance, its length of 2s approximates the scan acquisition times needed for a single multi-slice T2* weighted brain shot typical for functional MRI studies (Mórocz et al., 2012; Janoos et al., 2013). To reduce dimensionality and capture coherent spatial patterns, we grouped EEG channels based on the similarity of their time series. Pairwise distances between channels were computed, followed by hierarchical clustering to identify groups of similar channels. For the feature vector construction, we computed at each time step the mean signal intensity across predefined frequency bands within each cluster of grouped channels. This resulted in a concatenated vector representation summarizing the spatio-spectral profile of the EEG activity at each moment. For the identification of brain states, we applied k-means clustering to the resulting vectors to identify a set of prototypical spatio-spectral configurations, which we interpret as data-driven representations of transient brain states.
@@ 3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. We observe that a set of feature combinations of high importance include both eyes-open and eyes-closed conditions and the delta frequency band that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in C.
@@ 3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These two clinical measures arguably reflect disease related accumulation of neurological deficits.
@@
The plots in figure 3 imply a correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden over time accumulates in the CNS. Patients with symptoms other than pure motor symptoms show a subtle trend with EDSS computed on 26 such cases (middle graph B). Regarding disease duration, however, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
@@
Figure 3: Distinction of control subjects from MS patients through brain states prediction. Disease predictor values correspond to the rotated coordinates as seen in Fig. 1B using principal components of the point distribution with the predictor value considered as primary principal component. (A) The upper graph illustrates this separability for both cohorts plotted against EDSS measures obtained in MS. Patients are shown as red or green dots and controls as blue. Seven outlier patients are found on the left marked with green characters (a-g) for further discussion under Results. (B) The middle graph plots separability against EDSS measures in MS patients alone. Only patients having symptoms other then motor symptoms were included. For these 26 patients a linear regression is shown as gray line. (C) The lower graph plots separability against disease duration in years. Six patients (including d,f) are excluded for no data on disease duration. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and EDSS or disease duration. The open-source python module linregress from scipy served for calculation (Virtanen et al., 2020).
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into a new analytic method for processing EEG data using a space-neutral quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-neutral data-driven full-brain raw-data EEG analysis approach for MS.
@@ 4.3 Limitations
Third, the difference of mean age between the two cohorts amounts to nine years. We are, however, for the current retrospective analysis not aware of a correlation between age and disease severity (EDSS or symptoms) or disease duration. All subjects still belong to a young life stage where we expect task-less experiments to recruit similar brain functions and areas and so not to bias age-specifically the brain states predictor space.
Fourth, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. Hence this window size will be optimized in future iterations of this project.
Fifth, the current implementation of the BSD method is not suited for the topographic visualization of the 8 brain states, as it is in the case of the microstates method that follows a clear anatomical preponderance (Lehmann et al., 1987). Our current method focuses on a space-neutral approach for analyzing EEG data of patients with MS.
@@ 4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free space-neutral data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).

### v8

@@ Abstract
Objective: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
@@ 1 Introduction
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
@@ 2.3 Brain states detection (BSD)
In a first move a limited set of eight brain states are computed. Raw EEG signals from each channel are transformed into sequences of overlapping spectrograms using the Fast Fourier Transform (FFT) implementation from the cusignal.spectrogram library (Thompson and Nicely, 2021). The recordings are segmented into data epochs where each individual window, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz). While window size and number of states are varied as technical optimization against detection performance, its length of 2s approximates the scan acquisition times needed for a single multi-slice T2* weighted brain shot typical for functional MRI studies (Mórocz et al., 2012; Janoos et al., 2013). To reduce dimensionality and capture coherent spatial patterns, we grouped EEG channels based on the similarity of their time series. Pairwise distances between channels were computed, followed by hierarchical clustering to identify groups of similar channels. For the feature vector construction, we computed at each time step the mean signal intensity across predefined frequency bands within each cluster of grouped channels. This resulted in a concatenated vector representation summarizing the spectral profile of the EEG activity at each moment. For the identification of brain states, we applied k-means clustering to the resulting vectors to identify a set of prototypical channel-spectral configurations, which we interpret as data-driven representations of transient brain states.
@@ 3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. To reduce the risk of overfitting only the first 15 feature combinations of high importance including both eyes-open and eyes-closed conditions and the delta frequency band were selected that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in C.
@@ 3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as age, EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. Age itself does not indicate an observable effect on MS prediction in both cohorts (top graph A in figure 3). However the next two clinical measures arguably reflect disease related accumulation of neurological deficits. The next three plots in figure 3 imply a correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden over time accumulates in the CNS. Patients with symptoms other than pure motor symptoms show a subtle trend with EDSS computed on 26 such cases (third graph C). Regarding disease duration, however, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
@@
Figure 3: Distinction of control subjects from MS patients through brain states prediction. Disease predictor values correspond to the rotated coordinates as seen in Fig.1B using principal components of the point distribution with the predictor value considered as primary principal component. (A) The top graph illustrates age distribution as a function of the MS predictor scale. It indicates no observable effect of age on MS prediction. The control outliers on the right side (false positives in blue) do not have higher ages than the controls on the left side (true negatives). Conversely, the MS cases on the left side (false negatives in green) do not have a lower age distribution than the true positives on the right side. (B) The second graph illustrates this separability in the MS cohort plotted against EDSS measures. Patients are shown as red or green dots as in the top graph. The seven outliers on the left are marked with green characters (a-g) for further discussion under Results. (C) The third graph plots separability against EDSS measures only for MS patients having symptoms other then motor symptoms. For these 26 patients a linear regression is shown as gray line. (D) The bottom graph plots separability against disease duration in years. Six patients (including d,f) are excluded for no data on disease duration. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and EDSS or disease duration. The open-source python module linregress from scipy served for calculation (Virtanen et al., 2020).
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into a new analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free data-driven full-brain raw-data EEG analysis approach for MS.
@@ 4.3 Limitations
Third, the difference of mean age between the two cohorts amounts to nine years. We are, however, for the current retrospective analysis not aware of a correlation between age and disease severity (EDSS or symptoms) or disease duration. All subjects still belong to a life stage where we expect task-less experiments to recruit similar brain functions and areas and hence to not age-specifically bias the brain states predictor scale.
Fourth, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. Hence this window size will be optimized in future iterations of this project with a multi-STFT approach (Ali et al., 2022).
Fifth, the current implementation of the BSD method is not suited for the topographic visualization of the 8 brain states, as it is in the case of the microstates method that follows a clear anatomical preponderance (Lehmann et al., 1987). The spatial information was for code-technical and computational performance optimization reasons lost. Our current method focuses on a channel-spectral approach for analyzing EEG data of patients with MS.
@@ 4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).